COMETARY ICE ANALOGS: TPD AND FTIR STUDY OF CO_{2}/H_{2}O ICES

\underline{VICTOR J. HERRERO}, R. ESCRIBANO, B. MATÉ, O. GÁLVEZ, I.K. ORTEGA, B. MARTIN-LLORENTE, Instituto de Estructura de la Materia, CSIC, Madrid, Spain; P.J. GUTIÉRREZ, Instituto de Astrofísica de Andalucía, CSIC, P.O.Box 3004, 18080 Granada, Spain.

Frozen CO_{2} is, after water, one of the main components of cometary nuclei. The structure of CO_{2} ice in interaction with water ice depends on several factors, like deposition scheme, temperature or density. We present an investigation on ice mixtures of water and CO_{2} and their variation with temperature, based on temperature programmed desorption (TPD) and infrared spectroscopy measurements. Two slightly different CO_{2} structures are shown to exist, and their relative concentration is found to depend critically on the deposition sequence and temperature of the sample.